# MotionCtrl: A Unified and Flexible Motion Controller for Video Generation https://wzhouxiff.github.io/projects/MotionCtrl/

### v1

@@ 3.2.3 Training Strategy and Data Construction
Learning the camera motion control module (CMCM) requires a training dataset that contains video clips with captions, and camera poses, but not object movement trajectories. To this end, we opt to use the Realestate10K dataset [33], which, after removing invalid video links, offers 62,992 video clips accompanied by diverse camera poses. Nonetheless, employing Realestate10K presents two potential challenges: 1) the dataset’s limited diversity of scenes, primarily from real estate videos, potentially compromising the quality of the generated video; and 2) it lacks captions needed for T2V models.
@@
Regarding the first challenge, we specifically design the CMCM module to only train several extra MLP layers and the second self-attention module of the temporal transformers in LVDM while freezing all other parameters (as illustrated in Figure 2 (b).). The training of CMCM mainly focuses on learning global motion and seldom affects the content of the generated video. Therefore, the limited scene diversity in the Realestate10K dataset has a negligible effect on the generated quality of LVDM. This is substantiated by quantitative results presented in Table 2, where the FID [22] and FVD [24] metrics indicate that the video quality generated by our MotionCtrl is on par with the LVDM outcomes.
To address the second challenge, we utilize Blip2 [13], an image captioning algorithm, to generate captions for each video clip in Realestate10K. We extract frames at specific intervals—the first, quarter, half, three-quarters, and final frames of the video. We then use Blip2 to predict their captions. These captions are then concatenated to form a comprehensive description for each video clip. With these captions in place, we train the CMCM on Realestate10K, which in turn allows the LVDM to control camera motion effectively.
Learning the object motion control module (OMCM) requires a dataset comprising video clips with corresponding captions and object movement trajectories. Currently, there is an absence of large-scale datasets that combine video-text pairs with object trajectories. To address this, we employ ParticleSfM [31] to generate object movement trajectories using the WebVid dataset [3], which is widely used for T2V generation. ParticleSfM features a trajectory motion segmentation network capable of identifying trajectories associated with moving objects within a video, as depicted in Figure 3 (b), where the trajectories predominantly correspond to a moving person. Despite its effectiveness, ParticleSfM is not time-efficient, requiring approximately 2 minutes to process a 32-frame video. To balance time efficiency and data coverage, we opt to randomly select 32 frames from each WebVid video, with a frame skip interval $s\in[1,16]$, to precompute the object movement trajectories. This approach yields a total of 243,000 samples that fulfill the training requirements for the OMCM.
To enhance user experience and interaction, we utilize sparse trajectories to direct object motion instead of dense ones. During the training process, we randomly select $n\in[1,N]$ trajectories (where $N$ is the maximum number of trajectories for each video) from the synthesized trajectories. An example of the outcome can be seen in Figure 3 (c). However, these selected sparse trajectories can be too scattered for effective training. To address this issue, we apply a Gaussian filter to the sparse trajectories (Figure 3 (d)), inspired by DragNUWA [28]. During training, we initially train the OMCM using dense trajectories and then proceed to fine-tune it with sparse trajectories.
@@
In this training phase, we adopt the LVDM model that is already fine-tuned with CMCM. We only train the OMCM layers, while the entire base model and CMCM remains frozen. This strategy guarantees that OMCM adds the object motion control capabilities with a limited dataset while minimally impacting LVDM and CMCM. Upon completion of this training phase, giving both camera poses and trajectories allows for flexible customization of camera and object motion in the generated video.
@@
Figure 3: Trajectories for Object Motion Control. ParticleSfM [31] is employed to extract object movement trajectories from video clips, effectively disentangling object motion from camera-induced movement. To circumvent the issues of dense trajectories, which can encode object shapes and are challenging to design at inference, we train the OMCM using sparse trajectories sampled from the dense ones. These sparse trajectories, being too scattered for effective learning, are subsequently refined with a Gaussian filter.
4 Experiments
4.1 Experiment Settings
Implementation Details. Our MotionCtrl is based on the LVDM [9] and the trained weights are provided by VideoCraft1 [5]. It is trained on 16-frame sequences at a resolution of $256\times 256$. We set the maximum number of trajectories $N$ to 8. Both CMCM and OMCM are optimized using the Adam optimizer [12] with a batch size of 128 and a learning rate of $1{\times}10^{-4}$ on 8 NVIDIA Tesla V100 GPUs. The CMCM requires approximately 50,000 iterations to converge. OMCM is initially trained on dense trajectories for 20,000 iterations, followed by fine-tuning with sparse trajectories for an additional 20,000 iterations.
Evaluation Datasets. We independently evaluate the efficacy of our proposed MotionCtrl on camera motion control and object motion control. These assessments are conducted using their respective evaluation datasets. More details about the two evaluation datasets are provided in the supplementary materials.
Camera motion control evaluation dataset. MotionCtrl’s camera motion control is tested on a set of 8 basic camera pose sequences: pan left, pan right, pan up, pan down, zoom in, zoom out, anticlockwise rotation, and clockwise rotation. Additionally, we randomly select 20 complicated camera pose sequences from the Realestate10K test set [33]. For each camera pose sequence, we evaluate 10 different text prompts.
Object motion control evaluation dataset. For object motion control, we create 19 diverse trajectories, including straight lines, curves, shaking lines, and complex combinations of those. For each trajectory, we evaluate 12 different text prompts.
Evaluation Metrics. 1) The quality of the generated videos is evaluated using Fréchet Inception Distance (FID)[22], Fréchet Video Distance (FVD)[24], and CLIP Similarity (CLIPSIM[16], which measure the visual quality, temporal coherence, and semantic similarity to the input text, respectively. Denoted that the reference videos of FID and FVD are 1000 videos selected from WebVid [3]. 2) The efficacy of the camera motion control and object motion control is quantified by computing the Euclidean distance between the predicted and ground truth camera poses and object trajectories. The camera poses and object trajectories for the predicted videos are extracted using ParticleSfM [31]. We title these two metrics as CamMC and ObjMC, respectively.
@@ 4.1 Experiment Settings
Figure 4: Qualitative Comparisons on Camera and Object Motion Control. (a) Basic Poses: MotionCtrl and AnimateDiff[8] effectively execute zooms, but MotionCtrl can adjust to varying camera moving speeds. (b) Complex Poses: VideoComposer[25] uses Realestate10K’s raw video for motion vectors, capturing unintended shapes like doors, leading to unnatural results (refer to frame 12). MotionCtrl, however, produces a relatively natural video with motion that closely matches the camera poses. (c) Object Trajectory: Both VideoComposer and MotionCtrl can generate an object moving along a given trajectory (red curve), but MotionCtrl more precisely follows it in each frame, as indicated by green points.
4.2 Comparisons with State-of-the-Art Methods
@@
To validate the effectiveness of our MotionCtrl in controlling both camera and object motion, we compare it with two leading methods: AnimateDiff [8] and VideoComposer [25]. AnimateDiff employs 8 separate LoRA [11] models to control 8 basic camera motions in videos, such as panning and zooming, while VideoComposer manipulates video motion using motion vectors without differentiating between camera and object movements. Although DragNUWA [28] is relevant to our research, its code is not publicly available, precluding a direct comparison. Moreover, DragNUWA only learns motion control with the trajectories extracted from optical flow, which cannot fine-grainedly distinguish the movement between foreground objects and background, limiting its ability to precisely control camera and object motion.
@@
We compare our MotionCtrl with these methods in terms of camera motion and object motion control, and show the capability of our MotionCtrl in flexibly combining the control of camera motion and object motion in video generation. More comparisons and video comparisons are provided in the supplementary materials.
@@
Camera Motion Control. We assess camera motion control using basic poses and complex poses from the Realestate10K test set. AnimateDiff [8] is limited to basic camera poses, while VideoComposer [25] handles complex poses by extracting motion vectors from provided videos. The qualitative results are shown in Figure 4. For basic poses, both MotionCtrl and AnimateDiff can produce videos with forward camera movement, but MotionCtrl can generate camera motion with varying speeds, while AnimateDiff is nonadjustable. Regarding complex poses, where the camera first moves left front and then forward, VideoComposer can mimic the reference video’s camera motion using extracted motion vectors. However, the dense motion vectors inadvertently capture object shapes, such as a door’s outline in the reference video (frame 12), resulting in an unnatural-looking Eiffel Tower. MotionCtrl, guided by rotation and translation matrices, generates more natural-looking videos with camera motion close to the reference.
@@
Quantitative results in Table 1 show MotionCtrl’s superiority over AnimateDiff and VideoComposer for both basic and complex poses, as reflected by the CamMC score. Additionally, MotionCtrl achieves better text similarity and quality metrics, as measured by CLIPSIM, FID, and FVD.
Method AnimateDiff [8] VideoComposer [25] MotionCtrl
CamMC normal-$\downarrow$ (Basic Poses) 0.0548 - 0.0289
CamMC normal-$\downarrow$ (Complex Poses) - 0.1073 0.0840
ObjMC normal-$\downarrow$ - 34.7778 25.1198
CLIPSIM normal-$\uparrow$ 0.2144 0.2219 0.2324
FID normal-$\downarrow$ 157.73 134.97 130.29
FVD normal-$\downarrow$ 1815.88 1045.82 934.37
@@
Table 1: Quantitative Comparisons with SOTA Methods. Our MotionCtrl outperforms competing approaches in both camera and object motion control while also excelling at preserving text similarity and the quality of the video generation.
@@
Object Motion Control. We compare our MotionCtrl with VideoComposer for object motion control, where VideoComposer utilizes motion vectors extracted from trajectories. The qualitative results are shown in Figure 4 (c). The red curve illustrates the given trajectory, while the green points indicate the expected object locations in the corresponding frame. The visual comparison reveals that MotionCtrl can generate objects whose movements are closer to the given trajectories, whereas VideoComposer’s results deviate in certain frames, highlighting MotionCtrl’s superior object motion control capability. The quantitative results in Table 1 demonstrate that MotionCtrl achieves better object motion control than VideoComposer, as reflected by the improved ObjMC score.
@@
Combination of Camera Motion and Object Motion. MotionCtrl can not only control camera and object motion independently within a single video but also perform integrated control of both. As demonstrated in Figure 1 (b) and (c), when MotionCtrl is applied with only a trajectory, it primarily generates a swaying rose that follows this path. By further introducing zoom-out camera poses, both the rose and the background are animated in accordance with the specified trajectory and camera movements. Additional results can be found in the supplementary materials and on the project page.
4.3 Ablation Studies
In this section, we present experiments to validate the effectiveness of the delicate designs in MotionCtrl, including the integrated positioning of CMCM and LVDM, the training strategy for OMCM, and the sequence in which CMCM and OMCM are trained.
Integrated Position of Camera Motion Control Module (CMCM). We test implementing camera motion control by combining camera poses with the time embedding, spatial cross-attention, or spatial self-attention module in LVDM. Although such methods have succeeded in other types of controlling [29, 15, 14], such as sketch, depth, and so on, they fail to endow camera control capabilities to LVDM, as evidenced by the CamMC scores in Table 2, being close to LVDM without camera motion control. That is because these components primarily focus on spatial content generation, which is largely insensitive to the camera motion encoded in camera poses. Conversely, incorporating CMCM with LVDM’s temporal transformers significantly improves camera motion control, as indicated by a lower CamMC score of $0.0289$ in Table 2. Camera motion primarily causes global view transformations over time, and fusing camera poses into LVDM’s temporal blocks aligns with this property, enabling effective camera motion control during video generation.
Furthermore, our CMCM has a negligible impact on the quality of spatial content, with FID[22] and FVD[24] scores comparable to LVDM, even though it is trained on the relatively small Realestate10K dataset [33] with low diversity. In contrast, injecting camera poses through time embedding or spatial transformers compromises quality scores, as it disrupts spatial content creation.
The corresponding qualitative results are in the supplementary materials.
Method CamMC normal-$\downarrow$ CLIPSIM normal-$\uparrow$ FID normal-$\downarrow$ FVD normal-$\downarrow$
LVDM [9] 0.9010 0.2359 130.62 1007.63
@@ 4.3 Ablation Studies
Table 2: Ablation of Camera Motion Control. Our Camera Motion Control Module (CMCM), incorporated with the temporal transformers of LVDM [9], effectively controls camera motion and maintains LVDM’s video quality. Conversely, other variants fail to control the camera motion and may reduce the LVDM’s generative quality.
Dense Trajectories v.s. Sparse Trajectories. We first train OMCM with dense object movement trajectories extracted via ParticleSfM[31] and then fine-tune with sparse trajectories. We evaluate the effectiveness of this approach by comparing it with training OMCM solely on dense or sparse trajectories. Table 3 indicates that training exclusively with dense trajectories yields inferior outcomes. This can be attributed to discrepancies between the training and inference phases (since only sparse trajectories are available during inference). While training solely with sparse trajectories shows improvement over the dense-only approach, it still falls short of the hybrid method. Sparse trajectories alone provide insufficient information for OMCM’s learning. In contrast, dense trajectories offer richer information that accelerates learning, and subsequent fine-tuning with sparse trajectories allows OMCM to adjust to the sparsity encountered during inference.
The corresponding qualitative results are in the supplementary materials.
Method ObjMC normal-$\downarrow$ CLIPSIM normal-$\uparrow$ FID normal-$\downarrow$ FVD normal-$\downarrow$

### v2

@@ 3.2.3. Training Strategy and Data Construction
Learning the camera motion control module (CMCM)
@@
CMCM only requires a training dataset that contains video clips with annotations of captions and camera poses. Considering that Realestate10K (Zhou et al., 2018) contains over 60k videos with relatively clean annotations of camera poses, we take it as our training dataset of CMCM. However, employing Realestate10K in MotionCtrl presents two potential challenges: 1) The diversity of scenes is limited in Realestate10K, primarily from real estate videos, potentially compromising the quality of the generated video; and 2) it lacks captions needed for T2V models.
@@
Regarding the first challenge, we adopt an adapter-like control module (CMCM), with only several new added MLP layers and the second self-attention module of the temporal transformers in LVDM are trainable, and reserving the generation quality of LVDM by freezing most of its parameters. Since the temporal transformers are mainly focus on the learning of global motions, the limited scene diverisity of Realestate10K seldom affects the generation quality of LVDM. This is substantiated by quantitative results presented in Table 2, where the FID (Seitzer, 2020) and FVD (Unterthiner et al., 2018) metrics indicate that the video quality generated by our MotionCtrl is on par with the LVDM outcomes.
To address the second challenge, we adopt Blip2 (Li et al., 2023), an image captioning algorithm, to generate captions for each video clip in Realestate10K. Details are in the supplementary materials.
Learning the object motion control module (OMCM)
OMCM requires a dataset comprising video clips with captions and object movement trajectories, which is currently lacking in the community. To meet the requirement, we utilize ParticleSfM (Zhao et al., 2022) to synthesize object movement trajectories in WebVid (Bain et al., 2021). WebVid a large-scale video dataset equipped with captions and commonly used in the T2V generation task. Although ParticleSfM is a structure-from-motion system primarily, it incorporates a trajectory-based motion segmentation module utilized for filtering out dynamic trajectories that affect the production of camera trajectories in a dynamic scene. The dynamic trajectories attained by the motion segmentation module exactly fulfill the requirements of our MotionCtrl and we employ this module to synthesize moving object trajectories for about 243,000 videos in WebVid. An example is illustrated in Fig. 3 (b), where the trajectories predominantly correspond to a moving person. Synthesis details are in the supplementary.
To circumvent the necessity for users to provide dense trajectories as depicted in Fig. 3 (b), which may not be user-friendly, MotionCtrl is required to control the moving objects based on sparse (one or a few) trajectories provided by users. Consequently, our OMCM is trained with $n\in[1,N]$ trajectories (where $N$ represents the maximum number of trajectories for each video) randomly selected from the synthesized dense trajectories (as shown in Fig. 3 (c)). Nevertheless, these selected sparse trajectories tend to be too scattered for effective training. Drawing inspiration from DragNUWA (Yin et al., 2023a), we mitigate this issue by applying a Gaussian filter to the sparse trajectories (Fig. 3 (d)) and we initially train the OMCM using dense trajectories before fine-tune it using sparse trajectories.
@@
In this training phase, both LVDM and CMCM are well-trained and frozen, with only the OMCM is trained. This strategy guarantees that OMCM adds the object motion control capabilities with a limited dataset while minimally impacting LVDM and CMCM. Upon the completion of this training phase, giving both camera poses and object trajectories allows for flexible controlling the camera and object motion in the generated video.
@@
Figure 3. Trajectories for Object Motion Control. ParticleSfM (Zhao et al., 2022) is employed to extract object movement trajectories from video clips, effectively disentangling object motion from camera-induced movement. To circumvent the issues of dense trajectories, which can encode object shapes and are challenging to design at inference, we train the OMCM using sparse trajectories sampled from the dense ones. These sparse trajectories, being too scattered for effective learning, are subsequently refined with a Gaussian filter.
Refer to caption Refer to caption
(a) Camera motion control on basic poses (b) Camera motion control on relatively complex poses
@@
Figure 4. Qualitative Comparisons on Camera Motion Control. (a) Basic Poses: MotionCtrl and AnimateDiff(Guo et al., 2023) effectively execute zooms, but MotionCtrl can adjust to varying camera moving speeds. (b) Relatively Complex Poses: VideoComposer(Wang et al., 2023) uses Realestate10K’s raw video for motion vectors, capturing unintended shapes like doors, leading to unnatural results (refer to frame 12). MotionCtrl, however, produces a relatively natural video with motion that closely matches the camera poses.
4. Experiments
4.1. Experiment Settings
4.1.1. Implementation Details.
MotionCtrl is built upon the LVDM framework (He et al., 2022)/VideoCraft1 (Chen et al., 2023b), which is trained on 16-frame sequences at a resolution of $256\times 256$. It can be readily adapted to other video generation models with similar structures, such as AnimateDiff (Guo et al., 2023), adhering to the settings specific to each model. Additionally, the maximum number of trajectories $N$ is fixed at 8. Both CMCM and OMCM are optimized using the Adam optimizer (Kingma and Ba, 2014) with a batch size of 128 and a learning rate of $1e^{-4}$ across 8 NVIDIA Tesla V100 GPUs. The CMCM typically requires approximately 50,000 iterations to converge. Meanwhile, OMCM undergoes an initial training phase on dense trajectories for 20,000 iterations, followed by fine-tuning with sparse trajectories for an additional 20,000 iterations.
4.1.2. Evaluation Datasets.
(1) Camera motion control evaluation dataset encompasses two types of camera poses: basic camera poses (pan left, pan right, pan up, pan down, zoom in, zoom out, anticlockwise rotation, and clockwise rotation) and relatively complex camera poses111”Complex camera poses” in this work denotes camera movement beyond the basic camera poses. While basic camera poses involve movement in a single straight direction, complex camera poses contain movement in several directions. obtained from the test set of Realestate10K (Zhou et al., 2018) or synthesized using ParticleSfM (Zhao et al., 2022) on videos from WebVid (Bain et al., 2021) and HD-VILA (Xue et al., 2022). (2) Object motion control evaluation dataset consists of 283 samples constructed with diverse handcrafted trajectories and prompts. Further details regarding the construction of the evaluation datasets are provided in the supplementary materials.
4.1.3. Evaluation Metrics.
(1) The quality of the generated videos is evaluated using Fréchet Inception Distance (FID)(Seitzer, 2020), Fréchet Video Distance (FVD)(Unterthiner et al., 2018), and CLIP Similarity (CLIPSIM(Radford et al., 2021), which measure the visual quality, temporal coherence, and semantic similarity to the text, respectively. Denoted that the reference videos of FID and FVD are 1000 videos from WebVid (Bain et al., 2021). (2) The efficacy of the camera and object motion control is quantified by computing the Euclidean distance between the predicted and ground truth camera poses and object trajectories, respectively. The camera poses and object trajectories of the predicted videos are extracted using ParticleSfM (Zhao et al., 2022). We title these two metrics as CamMC and ObjMC, respectively. 3) We also conduct a user study for subjective quantitative evaluation, the details of which are provided in the supplementary materials due to space limitations.
Figure 5. Qualitative Comparisons on Object Motion Control. Both VideoComposer and MotionCtrl can generate an object moving along a given trajectory (red curve), but MotionCtrl more precisely follows it in each frame, as indicated by green points.
4.2. Comparisons with State-of-the-Art Methods
@@
To validate the effectiveness of our MotionCtrl in controlling both camera and object motion, we compare it with two leading methods: AnimateDiff (Guo et al., 2023) and VideoComposer (Wang et al., 2023). AnimateDiff employs 8 separate LoRA (Hu et al., 2021) models to control 8 basic camera motions in videos, such as panning and zooming, while VideoComposer manipulates video motion using motion vectors without differentiating between camera and object movements. Although DragNUWA (Yin et al., 2023a) is relevant to our research, its code is not publicly available, precluding a direct comparison. Moreover, DragNUWA only learns motion control with the trajectories extracted from optical flow, which cannot fine-grainedly distinguish the movement between foreground objects and background, limiting its ability to precisely control camera and object motion.
@@
We compare our MotionCtrl with these methods in terms of camera motion and object motion control, and show the capability of our MotionCtrl to flexibly combine the control of camera motion and object motion in video generation. More comparisons and video comparisons are provided in the supplementary materials.
4.2.1. Camera Motion Control.
@@
We assess camera motion control using basic poses and relatively complex poses. AnimateDiff (Guo et al., 2023) is limited to basic camera poses, while VideoComposer (Wang et al., 2023) handles complex poses by extracting motion vectors from provided videos. The qualitative results are shown in Fig. 4. For basic poses, both MotionCtrl and AnimateDiff can produce videos with forward camera movement, but MotionCtrl can generate camera motion with varying speeds, while AnimateDiff is nonadjustable. Regarding complex poses, where the camera first moves left front and then forward, VideoComposer can mimic the reference video’s camera motion using extracted motion vectors. However, the dense motion vectors inadvertently capture object shapes, the door’s outline in the reference video (frame 12), resulting in an unnatural-looking Eiffel Tower. MotionCtrl, guided by rotation and translation matrices, generates more natural-looking videos with camera motion close to the reference.
@@
Quantitative results in Table 1 show MotionCtrl’s superiority over AnimateDiff and VideoComposer for both basic and relatively complex poses, as reflected by the CamMC score. Additionally, MotionCtrl achieves better text similarity and quality metrics, as measured by CLIPSIM, FID, and FVD.
@@
Table 1. Quantitative Comparisons with AnimateDiff (Guo et al., 2023) and VideoComposer (Wang et al., 2023). Our MotionCtrl outperforms competing approaches in both camera and object motion control while also excelling at preserving text similarity and the quality of the video generation.
Method AnimateDiff VideoComposer MotionCtrl
CamMC $\downarrow$ (Basic Poses) 0.0548 - 0.0289
CamMC $\downarrow$ (Complex Poses) - 0.0950 0.0735
ObjMC $\downarrow$ - 36.8351 28.877
CLIPSIM $\uparrow$ 0.2144 0.2214 0.2319
FID $\downarrow$ 157.73 130.97 124.09
FVD $\downarrow$ 1815.88 1004.99 852.15
4.2.2. Object Motion Control.
@@
We compare our MotionCtrl with VideoComposer for object motion control, where VideoComposer utilizes motion vectors extracted from trajectories. The qualitative results are shown in Fig. 5. The red curve illustrates the given trajectory, while the green points indicate the expected object locations in the corresponding frame. The visual comparison reveals that MotionCtrl can generate objects whose movements are closer to the given trajectories, whereas VideoComposer’s results deviate in certain frames, highlighting MotionCtrl’s superior object motion control capability. The quantitative results in terms of ObjMC in Table 1 also demonstrate that MotionCtrl achieves better object motion control than VideoComposer.
4.2.3. Combination of Camera Motion and Object Motion.
@@
MotionCtrl can not only control camera and object motion independently within a single video but also perform integrated control of both. As demonstrated in Fig. 1 (b) and (c), when MotionCtrl is applied with only a trajectory, it primarily generates a swaying rose that follows this path. By further introducing zoom-out camera poses, both the rose and the background are animated in accordance with the specified trajectory and camera movements.
More results of MotionCtrl can be found in Fig. 8, supplementary materials, and the demo video.
Table 2. Ablation of Camera Motion Control. Our Camera Motion Control Module (CMCM), incorporated with the temporal transformers of LVDM (He et al., 2022), effectively controls camera motion and maintains LVDM’s video quality.
Method CamMC $\downarrow$ CLIPSIM $\uparrow$ FID $\downarrow$ FVD $\downarrow$
LVDM (He et al., 2022) 0.9010 0.2359 130.62 1007.63
@@
4.3. Ablation Studies
4.3.1. Integrated Position of Camera Motion Control Module (CMCM).
We test implementing camera motion control by combining camera poses with the time embedding, spatial cross-attention, or spatial self-attention module in LVDM. Although such methods have succeeded in other types of controlling (Zhang et al., 2023; Mou et al., 2024), such as sketch and depth, they fail to endow camera control capabilities to LVDM, as evidenced by the CamMC scores in Table 2 and visualized results in Fig. 6. Their CamMC scores are close to the original LVDM. That is because these components primarily focus on spatial content generation, which is insensitive to the camera motion encoded in camera poses. Conversely, incorporating CMCM with LVDM’s temporal transformers significantly improves camera motion control, as indicated by a lower CamMC score of $0.0289$ in Table 2. Camera motion primarily causes global view transformations over time, and fusing camera poses into LVDM’s temporal blocks aligns with this property, enabling effective camera motion control during video generation.
4.3.2. Dense Trajectories v.s. Sparse Trajectories.
OMCM is initially trained with dense object movement trajectories extracted via ParticleSfM (Zhao et al., 2022) and then fine-tune with sparse trajectories. We evaluate the effectiveness of this approach by comparing it with training OMCM solely on dense or sparse trajectories. Table 3 and Fig. 7 indicate that training exclusively with dense trajectories yields inferior outcomes, which is attributed to discrepancies between the training and inference phases (sparse trajectories are provided during inference). Though training solely with sparse trajectories shows improvement over the dense-only approach, it still falls short of the hybrid method, since sparse trajectories alone provide limited information. In contrast, dense trajectories offer richer information that accelerates learning, and subsequent fine-tuning with sparse trajectories allows OMCM to adjust to the sparsity encountered during inference.
Table 3. Ablation of Object Motion Control. The Object Motion Control Module (OMCM), when initially trained on dense object movement trajectories and subsequently fine-tuned with sparse trajectories, outperforms versions trained exclusively on either dense or sparse trajectories.
Method ObjMC $\downarrow$ CLIPSIM $\uparrow$ FID $\downarrow$ FVD $\downarrow$